OpenStudy (anonymous):

Help please Harry had $32. He spent all the money on buying 3 notebooks for $x each and 4 packs of index cards for $y each. If Harry had bought 5 notebooks and 5 packs of index cards, he would have run short of $18. A student concluded that the price of each notebook is $5 and the price of each pack of index cards is $1. Which statement best justifies whether the student's conclusion is correct or incorrect?

OpenStudy (anonymous):

The student's conclusion is incorrect because the solution to the system of equations 3x + 4y = 32 and 5x + 5y = 50 is (8, 2). The student's conclusion is incorrect because the solution to the system of equations 3x + 4y = 32 and 5x + 5y = 18 is (8, 2). The student's conclusion is correct because the solution to the system of equations 3x + 4y = 32 and 5x + 5y = 18 is (5, 1). The student's conclusion is correct because the solution to the system of equations 3x – 4y = 32 and 5x – 5y = 50 is (5, 1).

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

Hello @sdkfhldjk and Welcome to OpenStudy! :) for this problem, we should first try to convert the information they gave us into equations the idea is that you would multiply the number of object by that object's price to get the cost of the total objects so 3*x + 4*y = 32 because 3 is the number of notebooks and x is the price of the notebooks - so they get multiplied together so you see how I got that equation now? :)

OpenStudy (anonymous):

I sort of get it

OpenStudy (anonymous):

How would I do the second equation then? @jigglypuff314

jigglypuff314 (jigglypuff314):

the second equation would be 5*x + 5*y = (30+18) which can simplify into the equation that you need ^_^

OpenStudy (anonymous):

Thank you! @jigglypuff314